# Optimal low-rank approximations of posteriors for linear Gaussian inverse problems on Hilbert spaces

### v1

@@ B.2 Proofs of Section 4
For an arbitrary fixed $i$, if $f^{\prime}(\gamma_{i})\not=0$, then this implies that $U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$, as otherwise there exists $\varphi\in\mathbb{R}^{r}$ such that $\langle\varphi,U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}\rangle_{\mathbb{R}^{r}}\not=0$ and the choice $V=\mathcal{C}_{\textup{pos}}^{1/2}e_{i}\otimes\varphi$ furnishes a contradiction. Since $f\in\mathscr{F}$, it follows from Lemma 4.1(i) that $f^{\prime}(\gamma_{i})=0$ if and only if $\gamma_{i}=0$. Hence, $\gamma_{i}U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$ for all $i$.
By (17), (3c) and Lemma A.13(i),
$\mathcal{C}_{\textup{pos}}^{1/2}(UU^{*}-H)\mathcal{C}_{\textup{pos}}^{1/2}=\mathcal{C}_{\textup{pos}}^{1/2}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}-\mathcal{C}_{\textup{pos}}^{-1})\mathcal{C}_{\textup{pos}}^{1/2}\subset g(U).$
As the domain of $\mathcal{C}_{\textup{pos}}^{1/2}(UU^{*}-H)\mathcal{C}_{\textup{pos}}^{1/2}$ is in fact all of $\mathcal{H}$, by continuity of $g(U)$ it holds that
$\mathcal{C}_{\textup{pos}}^{1/2}(UU^{*}-H)\mathcal{C}_{\textup{pos}}^{1/2}=g(U)=\sum_{i}\gamma e_{i}\otimes e_{i}.$
Multiplying this equation from the right and using $\gamma U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$ for all $i$,
$\mathcal{C}_{\textup{pos}}^{1/2}(UU^{*}-H)\mathcal{C}_{\textup{pos}}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$ $=\left(\sum_{i}\lambda_{i}e_{i}\otimes e_{i}\right)\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$
$=\left(\sum_{i}e_{i}\otimes(\lambda_{i}U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i})\right)U^{*}\mathcal{C}_{\textup{pos}}^{1/2}=0.$
Thus, if $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ is a stationary point of $J_{f}$, then
@@
By injectivity of $\mathcal{C}_{\textup{pos}}^{1/2}$ and Lemma A.1, we have $\operatorname{rank}\left({\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}}\right)=\operatorname{rank}\left({UU^{*}}\right)=\operatorname{rank}\left({U}\right)$. Hence $\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$ is a non-negative and self-adjoint operator of rank at most $r$. In particular, it has $k$ many nonzero eigenvalues for some $k\leq r$. It follows from (45) that these $k$ eigenpairs $(\lambda,v)\in\mathbb{R}\setminus\{0\}\times\mathcal{H}$ of $\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$ are also eigenpairs of $\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}$. By Lemma 3.6, there exists an increasing sequence $(\lambda_{i})_{i}\in\ell^{2}((-1,0))$ with exactly $\operatorname{rank}\left({H}\right)$ nonzero entries and an ONB $(w_{i})_{i}$ of $\mathcal{H}$ with $(w_{i})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ such that
$\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}=\sum_{i}(-\lambda_{i})(\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i})\otimes(\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}).$
It follows that there exists a set of $k$ distinct indices $\mathcal{I}\coloneqq\{i_{1},\ldots,i_{k}\}\subset\{1,\ldots,\operatorname{rank}\left({H}\right)\}$ such that
$\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$ $=\sum_{j=1}^{k}(-\lambda_{i_{j}})\sqrt{1+\lambda_{i_{j}}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i_{j}}\otimes\sqrt{1+\lambda_{i_{j}}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i_{j}}$
$=\sum_{j=1}^{k}(-\lambda_{i_{j}})\sqrt{1+\lambda_{i_{j}}}^{-1}\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i_{j}}\otimes\sqrt{1+\lambda_{i_{j}}}^{-1}\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i_{j}},$
where we use (10c) for the second equality. If $U$ is a stationary point, then $U\in\mathcal{B}(\mathbb{R}^{r},H)$ satisfies the previous equation, and hence
$U=\sum_{j=1}^{k}\sqrt{\frac{-\lambda_{i_{j}}}{1+\lambda_{i_{j}}}}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i_{j}}\otimes\psi_{j}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{V}),$
for some orthonormal sequence $(\psi_{1},\ldots,\psi_{k})\subset\mathbb{R}^{r}$, with,
@@
$\mathcal{V}:$ $=\operatorname{span}{\left(\mathcal{C}_{\textup{pr}}^{-1/2}w_{i},i=1,\ldots,\operatorname{rank}\left({H}\right)\right)}\subset\mathcal{H}.$
@@
By (10c) and the relation $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}$ given in Lemma 3.6, it holds for all $i$ that $v_{i}=\sqrt{1+\lambda_{i}}^{-1}\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}$. Thus, by (18), (10b) and the expansion of $\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$ given above,
$g(U)$ $=(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}-I+\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$
$=\sum_{i}\lambda_{i}\frac{\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}}{\sqrt{1+\lambda_{i}}}\otimes\frac{\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}}{\sqrt{1+\lambda_{i}}}+\sum_{j=1}^{k}(-\lambda_{i_{j}})\frac{\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i_{j}}}{\sqrt{1+\lambda_{i_{j}}}}\otimes\frac{\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i_{j}}}{\sqrt{1+\lambda_{i_{j}}}}.$
@@
Thus, if $U$ is a stationary point of $J_{f}$, then the eigenvalues of $g(U)$ form the set $\{0\}\cup\{\lambda_{i}:\ i\not\in\mathcal{I}\}\subset(-1,0]$. Since $f\in\mathscr{F}$, it holds that $f(0)=0$, and it then follows from (19) that $J_{f}(U)=\sum_{i\not\in\mathcal{I}}f(\lambda_{i})$. Furthermore, $f$ is decreasing on $(-1,0]$, from which we conclude that for $\hat{U}$ to be a minimizer of $J_{f}$ among the stationary points of $J_{f}$, the $(\lambda_{i})_{i\in I}$ should be chosen as the $r$ most negative elements of $(\lambda_{i})_{i}$. That is, we take $\mathcal{I}=\mathcal{I}^{\textup{opt}}$ where $\mathcal{I}^{\textup{opt}}$ is any index set for which $\{\lambda_{i}:\ i\in\mathcal{I}^{\textup{opt}}\}=\{\lambda_{1},\ldots,\lambda_{r}\}$. The set $\mathcal{I}^{\textup{opt}}$ is uniquely defined if and only if $\lambda_{r}<\lambda_{r+1}$, in which case $\mathcal{I}^{\textup{opt}}=\{1,\ldots,r\}$. For such $\hat{U}$ we have
@@
and $J_{f}(\hat{U})=\sum_{i>r}f(\lambda_{i})$. If $\lambda_{r}=0$, then $\hat{U}$ and $\hat{U}\hat{U}^{*}$ are uniquely defined. If $\lambda_{i}\not=0$, then $\hat{U}$ and $\hat{U}\hat{U}^{*}$ are uniquely defined if and only if $\mathcal{I}^{\textup{opt}}$ is uniquely defined. Hence $\hat{U}\hat{U}^{*}$ is unique if and only if either $\lambda_{r}=0$ or $\lambda_{r}<\lambda_{r+1}$.
@@
By Lemma 4.16, $\hat{U}$ minimizes $J_{f}$ over $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$. Furthermore, if $\tilde{U}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ with $\operatorname{ran}{\tilde{U}}\not\subset\mathcal{V}$, then $\tilde{U}$ is not a stationary point of $J_{f}$, because a necessary condition for $\tilde{U}$ to be a stationary point of $J_{f}$ is that $\operatorname{ran}{\tilde{U}}\subset\mathcal{V}$. On the other hand, since $\hat{U}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{V})\subset\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{\tilde{U}})$, since $\tilde{U}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{\tilde{U}})$ and since $J_{f}$ is coercive over $\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{\tilde{U}})$ by Lemma 4.16, it follows that $J_{f}(\tilde{U})>J_{f}(\hat{U})$. Thus, $\hat{U}$ is a global minimiser of $J_{f}$. Since $J_{f}(\hat{U})=\mathcal{L}_{f}(\mathcal{C}_{\textup{pos}},(\mathcal{C}_{\textup{pr}}^{-1}+\hat{U}\hat{U}^{*})^{-1})$ by (19), it follows that the operator $\mathcal{P}^{\textup{opt}}_{r}$ in (20) solves 4.4.

### v3

@@ B.2 Proofs of Section 4
Since $f\in\mathscr{F}$, it follows from Lemma 4.1(i) that $f^{\prime}(\gamma_{i})=0$ if and only if $\gamma_{i}=0$. For an arbitrary fixed $j$, if $\gamma_{j}\not=0$, this implies that $U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{j}=0$ for a stationary point $U$, as otherwise there exists $\varphi\in\mathbb{R}^{r}$ such that $\langle\varphi,U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{j}\rangle_{\mathbb{R}^{r}}\not=0$ and the choice $V=\mathcal{C}_{\textup{pos}}^{-1/2}e_{j}\otimes\varphi$ furnishes a contradiction with $U$ being stationary. Indeed, in this case,
$\sum_{i}f^{\prime}(\gamma_{i})\langle\mathcal{C}_{\textup{pos}}^{1/2}e_{i},(\mathcal{C}_{\textup{pos}}^{-1/2}e_{j}\otimes\varphi)U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}\rangle$ $=\sum_{i}f^{\prime}(\gamma_{i})\langle\mathcal{C}_{\textup{pos}}^{1/2}e_{i},\mathcal{C}_{\textup{pos}}^{-1/2}e_{j}\rangle\langle U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i},\varphi\rangle$
=$f^{\prime}(\gamma_{j})\langle\varphi,U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{j}\rangle\not=0.$
Hence, if $U$ is a stationary point, then $\gamma_{i}U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$ for all $i$. Conversely, if $\gamma_{i}U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$ for all $i$, then for each $i$ it holds that $U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$ or $\gamma_{i}=0$, showing that $f^{\prime}(\gamma_{i})\langle\mathcal{C}_{\textup{pos}}^{1/2},VU^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}\rangle=0$ for all $i$ and all $V$. Hence $U$ is stationary if and only if $\gamma_{i}U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$ for all $i$.
Multiplying $g(U)=\sum_{i}\gamma_{i}e_{i}\otimes e_{i}$ from the right by $\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$ and using (19),
$\mathcal{C}_{\textup{pos}}^{1/2}(UU^{*}-H)\mathcal{C}_{\textup{pos}}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$ $=\left(\sum_{i}\gamma_{i}e_{i}\otimes e_{i}\right)\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}=\left(\sum_{i}e_{i}\otimes(\gamma_{i}U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i})\right)U^{*}\mathcal{C}_{\textup{pos}}^{1/2},$
where the second equation follows since $(u\otimes v)AA^{*}w=u\langle AA^{*}w,v\rangle=u\langle A^{*}w,A^{*}v\rangle=(u\otimes(A^{*}v))A^{*}w$ for suitable $u$, $v$, $w$, and $A$. Thus, if $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ is a stationary point of $J_{f}$, then
@@
Conversely, if (33) holds, then $\sum_{i}\langle\gamma_{i}U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i},U^{*}\mathcal{C}_{\textup{pos}}^{1/2}h\rangle e_{i}=0$ for all $h\in\mathcal{H}$, because (33) is equivalent to $\mathcal{C}_{\textup{pos}}^{1/2}(UU^{*}-H)\mathcal{C}_{\textup{pos}}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}=0$, and
$\mathcal{C}_{\textup{pos}}^{1/2}(UU^{*}-H)\mathcal{C}_{\textup{pos}}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}h=\left(\sum_{i}e_{i}\otimes(\gamma_{i}U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i})\right)U^{*}\mathcal{C}_{\textup{pos}}^{1/2}h=\sum_{i}\langle\gamma_{i}U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i},U^{*}\mathcal{C}_{\textup{pos}}^{1/2}h\rangle e_{i}.$
Since $(e_{i})_{i}$ is an ONB, this implies $\gamma_{i}\langle\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i},h\rangle=0$ for all $h\in\mathcal{H}$ and for all $i$. If $\gamma_{i}\not=0$ for some $i$, then taking $h=\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}$ we get $\lVert\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}\rVert=0$. Therefore, $\mathcal{C}_{\textup{pos}}^{1/2}e_{i}\in\ker{\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}}=\ker{UU^{*}}=\ker{U^{*}}$ by injectivity of $\mathcal{C}_{\textup{pos}}^{1/2}$ and Lemma A.7, showing $\gamma_{i}U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$ for all $i$. Thus, $U$ satisfies (33) if and only if $U$ is a stationary point.
@@
By injectivity of $\mathcal{C}_{\textup{pos}}^{1/2}$ and Lemma A.8, we have $\operatorname{rank}\left({\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}}\right)=\operatorname{rank}\left({\mathcal{C}_{\textup{pr}}^{1/2}U(\mathcal{C}_{\textup{pr}}^{1/2}U)^{*}}\right)=\operatorname{rank}\left({\mathcal{C}_{\textup{pos}}^{1/2}U}\right)=\operatorname{rank}\left({U}\right)$. Hence $\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$ is a non-negative and self-adjoint operator of rank at most $r$. In particular, it has $k$ many nonzero eigenvalues for some $k\leq r$. Suppose $U$ satisfies (33), so that the corresponding $k$ eigenpairs are also eigenpairs of $\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}$. By Proposition 3.7, there exists a nondecreasing sequence $(\lambda_{i})_{i}\in\ell^{2}((-1,0])$ with exactly $\operatorname{rank}\left({H}\right)$ nonzero entries and ONBs $(w_{i})_{i}$ and $(v_{i})_{i}$ of $\mathcal{H}$ with $w_{i},v_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ and $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}$ for all $i$, such that $\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}=\sum_{i}(-\lambda_{i})v_{i}\otimes v_{i}$. It follows that there exists a set of $k$ distinct indices $\{i_{1},\ldots,i_{k}\}\subset\{1,\ldots,\operatorname{rank}\left({H}\right)\}$ such that
$\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$ $=\sum_{j=1}^{k}(-\lambda_{i_{j}})v_{i_{j}}\otimes v_{i_{j}}.$ (34)
Conversely, if $U$ satisfies (34) for some distinct indices $\{i_{1},\ldots,i_{k}\}\subset\{1,\ldots,\operatorname{rank}\left({H}\right)\}$, then by using (34) and $\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}=\sum_{i}(-\lambda_{i})v_{i}\otimes v_{i}$ and direct computation, $U$ also satisfies (33). Thus, $U$ is a stationary point if and only if there exists an index set $\mathcal{I}\subset\{1,\ldots,\operatorname{rank}\left({H}\right)\}$ of cardinality at most $r$ and containing distinct indices such that
$UU^{*}$ $=\sum_{j\in\mathcal{I}}(-\lambda_{j})\mathcal{C}_{\textup{pos}}^{-1/2}v_{j}\otimes\mathcal{C}_{\textup{pos}}^{-1/2}v_{j}.$
In particular, by Lemma A.8, the stationary points $U$ of $J_{f}$ satisfy $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$, where,
@@
$\mathcal{V}$ $\coloneqq\operatorname{span}{\left(\mathcal{C}_{\textup{pos}}^{-1/2}v_{i},i=1,\ldots,\operatorname{rank}\left({H}\right)\right)}\subset\mathcal{H}.$
@@
Since $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}$ for all $i$, it holds by (10c) that $v_{i}=\sqrt{1+\lambda_{i}}^{-1}\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}$ for all $i$. Thus, if $U$ is a stationary point such that the above expression of $UU^{*}$ holds for the index set $\mathcal{I}$, then we have by (19) and (10b)
$g(U)$ $=\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}-\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}$
$=\sum_{j\in\mathcal{I}}(-\lambda_{j})\frac{\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{j}}{\sqrt{1+\lambda_{j}}}\otimes\frac{\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{j}}{\sqrt{1+\lambda_{j}}}+\sum_{i}\lambda_{i}\frac{\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}}{\sqrt{1+\lambda_{i}}}\otimes\frac{\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}}{\sqrt{1+\lambda_{i}}},$
@@
and hence the eigenvalues of $g(U)$ form the set $\{0\}\cup\{\lambda_{i}:\ i\not\in\mathcal{I}\}\subset(-1,0]$. Since $f\in\mathscr{F}$, it holds that $f(0)=0$, and it then follows from (20) that $J_{f}(U)=\sum_{i\not\in\mathcal{I}}f(\lambda_{i})$. Furthermore, $f$ is decreasing on $(-1,0]$. Let $\hat{U}$ be a stationary point corresponding to the index set $\mathcal{I}$. Then $\hat{U}$ minimises $J_{f}$ among its stationary points if and only if the sequence $(\lambda_{i})_{i\in\mathcal{I}}$ contains the $r$ most negative elements of $(\lambda_{i})_{i}$. In turn, this is the case if and only if $\mathcal{I}=\mathcal{I}^{\textup{opt}}$ where $\mathcal{I}^{\textup{opt}}$ is any index set for which $\{\lambda_{i}:\ i\in\mathcal{I}^{\textup{opt}}\}=\{\lambda_{1},\ldots,\lambda_{r}\}$. The set $\mathcal{I}^{\textup{opt}}$ is uniquely defined if and only if $\lambda_{r}<\lambda_{r+1}$, in which case $\mathcal{I}^{\textup{opt}}=\{1,\ldots,r\}$. Thus, using once more $v_{i}=\sqrt{1+\lambda_{i}}^{-1}\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}$, $\hat{U}$ satisfies
@@
and $J_{f}(\hat{U})=\sum_{i>r}f(\lambda_{i})$. Now, $\hat{U}\hat{U}^{*}$ is uniquely defined if and only if either $\lambda_{r+1}=0$ or $\mathcal{I}^{\textup{opt}}$ is uniquely defined. Hence $\hat{U}\hat{U}^{*}$ is unique if and only if either $\lambda_{r+1}=0$ or $\lambda_{r}<\lambda_{r+1}$.
@@
By Lemma 4.19, $\hat{U}$ minimizes $J_{f}$ over $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$ for the space $\mathcal{V}$ defined above. Furthermore, if $\tilde{U}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ with $\operatorname{ran}{\tilde{U}}\not\subset\mathcal{V}$, then $\tilde{U}$ is not a stationary point of $J_{f}$, because a necessary condition for $\tilde{U}$ to be a stationary point of $J_{f}$ is that $\operatorname{ran}{\tilde{U}}\subset\mathcal{V}$. On the other hand, since $\hat{U}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{V})\subset\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{\tilde{U}})$, since $\tilde{U}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{\tilde{U}})$ and since $J_{f}$ is coercive over $\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{\tilde{U}})$ by Lemma 4.19, it follows that $J_{f}(\tilde{U})>J_{f}(\hat{U})$. Thus, $\hat{U}$ is a global minimiser of $J_{f}$. Since $J_{f}(\hat{U})=\mathcal{L}_{f}(\mathcal{C}_{\textup{pos}}\|(\mathcal{C}_{\textup{pr}}^{-1}+\hat{U}\hat{U}^{*})^{-1})$ by (17), it follows that the operator $\mathcal{P}^{\textup{opt}}_{r}$ in (21) solves Problem 4.4.